# Offline Learning and Forgetting for Reasoning with Large Language Models

### v3

@@ Abstract
Leveraging inference-time search in large language models has proven effective in further enhancing a trained model’s capability to solve complex mathematical and reasoning problems. However, this approach significantly increases computational costs and inference time, as the model must generate and evaluate multiple candidate solutions to identify a viable reasoning path. To address this, we propose an effective approach that integrates search capabilities directly into the model by fine-tuning it on unpaired successful (learning) and failed reasoning paths (forgetting) derived from diverse search methods. A key challenge we identify is that naive fine-tuning can degrade the model’s search capability; we show this can be mitigated with a smaller learning rate. Extensive experiments on the challenging Game-of-24 and Countdown reasoning benchmarks show that, replacing CoT-generated data with search-generated data for offline fine-tuning improves success rates by around 23% over inference-time search baselines, while reducing inference time by 180. On top of this, our learning and forgetting objective consistently outperforms both supervised fine-tuning and preference-based methods.111Code is available at https://github.com/twni2016/llm-reasoning-uft.
@@ 1 Introduction
Chain-of-Thought (CoT) and inference-time search (search) have been proposed to enhance LLM generalization to reasoning problems at test time. Specifically, given a task description and input , CoT (Wei et al., 2022) outputs a single reasoning path via greedy decoding, or samples multiple paths to compute average performance. In contrast, search-based methods (Lightman et al., 2023; Yao et al., 2023a; Snell et al., 2024) explicitly generate intermediate reasoning (possibly with multiple passes) before extracting via a reward model or self-evaluation. Despite inference-time search’s effectiveness, it comes with greater computational costs (e.g., Fig. 1 and Chen et al. (2024b; a)).
@@
In this paper, we propose a straightforward method consisting of three phases designed to enhance the reasoning capabilities of large language models: data generation, fine-tuning, and evaluation. First, we collect heterogeneous reasoning data from both CoT and the search-based methods, and convert the data into a unified CoT-style path format. Next, we fine-tune the base LLM by learning successful paths with the supervised fine-tuning (SFT) and forgetting unpaired failed paths with the unlikelihood loss. We refer to this approach as unlikelihood fine-tuning (UFT). Finally, we evaluate the fine-tuned model on test sets using both CoT and inference-time search, with a primary focus on efficient CoT inference. We demonstrate the effectiveness of UFT on two math reasoning domains: Game-of-24 and Countdown, using Qwen2.5-Math 1.5B and 7B (Yang et al., 2024). We summarize our key findings below:
@@ 2 Related Work
Inference-time search has shown remarkable success across reasoning (Wang et al., 2022; Stechly et al., 2024; Hao et al., 2024; Snell et al., 2024) and planning tasks (Valmeekam et al., 2023; Zheng et al., 2024; Bohnet et al., 2024). Approaches include best-of-n sampling (Nakano et al., 2021; Wang et al., 2022), multi-agent debates (Du et al., 2023), and iterative correction via predefined rules (Bai et al., 2022) or self-refinement (Shinn et al., 2023; Madaan et al., 2024). A key direction, which this paper focuses on, integrates classic search algorithms with LLMs, employing BFS (Yao et al., 2023a; Xie et al., 2023), DFS (Qin et al., 2023), A* (Zhuang et al., 2023; Meng et al., 2024a; Koh et al., 2024) and MCTS (Hao et al., 2023; Zhou et al., 2023; Zhao et al., 2024; Xie et al., 2024; Gao et al., 2024). While effective, inference-time search methods remain computationally expensive: self-refinement requires multiple expensive passes, multi-agent approaches demand substantial GPU memory, and search trees grow exponentially with depth (Chen et al., 2024b).
@@
Synthetic data, generated by classic algorithms or LLMs, is widely used for SFT in reasoning tasks to mitigate the scarcity of human-annotated datasets (Liu et al., 2024). However, despite its abundance, synthetic responses may fail to solve the task. To address this, most prior work distill reasoning knowledge to policy (Hinton et al., 2015) by selecting only correct responses. Policy distillation strategies can be categorized based on whether the source and target data originate from CoT or search algorithms. CoT-to-CoT distillation learns CoT reasoning from correct CoT paths filtered by rewards, widely used in practice (Zelikman et al., 2022; 2024; Uesato et al., 2022; Yuan et al., 2023b; Gulcehre et al., 2023; Singh et al., 2023). Search-to-search distillation either learns a meta-policy by imitating entire search traces (Yang et al., 2022; Gandhi et al., 2024; Lehnert et al., 2024) or refines the proposal policy in tree search by imitating extracted paths (Feng et al., 2023; Tian et al., 2024; Zhang et al., 2024a). Notably, Zhang et al. (2024c) performs search-to-CoT distillation using preference pairs from Tree-of-Thought without a verifier. In comparison, we focus on CoT, search-to-CoT distillation from diverse reasoners and distill unpaired positive and negative data to the policy, labeled by a rule-based verifier.
@@
Negative data has been extensively studied in LLM safety, where the goal is to unlearn harmful content using objectives such as gradient ascent and unlikelihood training (Welleck et al., 2019; Keskar et al., 2019; Yao et al., 2023b; Zhang et al., 2024b). In preference optimization, negative examples are used to increase the relative likelihood of preferred responses in paired data , commonly used in alignment (Rafailov et al., 2023; Yuan et al., 2023a; Zhao et al., 2023; Hong et al., 2024; Xu et al., 2024; Meng et al., 2024b) and reasoning (Pal et al., 2024; Pang et al., 2024; Zhang et al., 2024c; Setlur et al., 2024). However, this paired format discards unpaired positives or negatives, reducing data efficiency. KTO (Ethayarajh et al., 2024) addresses the limitation of relative likelihood (Tuan & Wang, 2024) by learning from unpaired data, although it requires a reference model. In LLM reasoning, negative data is also used to train a reward model (Cobbe et al., 2021; Uesato et al., 2022; Lightman et al., 2023; Feng et al., 2023; Zhang et al., 2024a; Hosseini et al., 2024) for response re-ranking and search during inference. However, this does not inherently improve the base model as a policy. Instead of requiring preference pairs or reference models, we adopt unlikelihood training (Welleck et al., 2019) to directly forget failed paths as an auxiliary loss.
@@ 4.2 Unlikelihood Fine-Tuning on Unpaired Correct and Failed Paths
While prior work typically fine-tune LLMs on correct CoT-generated paths (Yuan et al., 2023b; Singh et al., 2023) or on preference pairs of correct and failed responses (Pang et al., 2024; Zhang et al., 2024c), we extend these approaches in two directions: (1) incorporating diverse reasoners beyond CoT to augment the training data, (2) fine-tuning the model to avoid and forget failed reasoning patterns using only failed examples. Unlike most preference-based methods, our approach leverages unpaired positive or negative data, without requiring success-failure pairs for the same question.
@@
This objective explicitly reduces the probability of incorrect paths under the model. Unlike gradient ascent (GA) unlearning (Yao et al., 2023b)444The GA objective is unstable to optimize (Zhang et al., 2024b; d), as it has an unbounded optimum () and a divergent gradient near the optimum. , UL objective has a smooth optimization landscape, with gradient:
@@
Since correct paths provide explicit solutions while wrong paths only rule out alternatives, we treat UL as an auxiliary loss by setting close to 0. Moreover, as correct and failed paths may be drawn from the same search trees, they often share prefixes, introducing conflicts when is large, as also noted in preference-based methods (Zhang et al., 2024c). For example, suppose a prefix sequence () appears in both and with equal frequency. The gradient of Eq. 5 w.r.t. this prefix is , which has a stationary point at . Thus, sets the desired probability of the shared prefix, further motivating the need for a small to avoid conflicting objectives. Unlike prior work (Zhang et al., 2024c; Setlur et al., 2024) that identify high-credit steps with more computation, we sidestep this with an auxiliary loss without preference pairs.
@@ 5 Experiments
We evaluate our approach on two mathematical reasoning tasks.
@@ 5.2 Boosting Chain-of-Thought Reasoning
Incoporating high-quality training data is decisive for CoT reasoning.
@@
Lastly, we analyze the effect of forgetting loss to SFT. First, we observe that UFT consistently improves over SFT in most scenarios: 14 out of 18 times for Q1.5B and 16 out of 18 for Q7B. In terms of magnitude, UFT yields an average improvement of 1.3% (with a maximum of 4%) in 3(b), and an average of 1.5% (with a maximum of 7%) in 2(b). In contrast, CPO-SimPO shows mixed results. While it improves over SFT with an average gain of 1.2% and maximum of 5% in 3(b), it consistently underperforms SFT in all scenarios in 2(b), with an average drop of 3.5%.
We believe several factors contribute to the contrasting behavior of CPO-SimPO relative to SFT (and UFT) when fine-tuned on Countdown versus Game-of-24. First, as shown in Appendix Table 5, the paired preference requirement in CPO-SimPO results in the exclusion of 0.1% to 5.8% of correct data on Countdown, while no such filtering occurs for Game-of-24. Since correct data determines data quality, this reduction likely contributes to CPO-SimPO’s weaker performance on Countdown. Second, the exclusion of unpaired failed data, along with the challenge of balancing preference and NLL objectives, may have a greater impact on the more difficult Countdown task.
@@ 5.3 Mitigating Forgetting in Search-based Reasoning
6 Conclusion
We propose a simple and effective method for mathematical reasoning, addressing three underexplored challenges: (1) augmenting training data by unifying mixed data formats from diverse reasoners into a CoT-style format, (2) leveraging unpaired correct and failed paths via unlikelihood fine-tuning, and (3) mitigating loss of search capability using a small learning rate. Experiments on Countdown and Game-of-24 show that data quality, unlikelihood loss, and learning rate are key to balancing CoT efficiency and search capability.
@@ A.2 Preference-based (Reference-free) Baselines: SimPO and CPO-SimPO
Appendix B Additional Results
@@ B.1 Training Dataset Statistics
B.2 Results on Qwen2.5-Math 7B
B.3 Results on Qwen2.5-Math 1.5B
Although not the primary focus of this work, we provide fine-tuning results based on Qwen2.5-Math 1.5B in Fig. 6 and Fig. 7 for completeness. Given its weak search-based reasoning performance as a base model (e.g., success rate 2% in Countdown) shown in Table 6, the effects of learning rate, CoT data, and unlikelihood loss are minimal, likely due to its poor initial capability. This suggests that the observed benefits depend on the model size.

### v4

@@ Abstract
Leveraging inference-time search in large language models has proven effective in further enhancing a trained model’s capability to solve complex mathematical and reasoning problems. However, this approach significantly increases computational costs and inference time, as the model must generate and evaluate multiple candidate solutions to identify a viable reasoning path. To address this, we propose an effective approach that integrates search capabilities directly into the model by fine-tuning it on unpaired successful (learning) and failed reasoning paths (forgetting) derived from diverse search methods. A key challenge we identify is that naive fine-tuning can degrade the model’s search capability; we show this can be mitigated with a smaller learning rate. Extensive experiments on the challenging Game-of-24 and Countdown arithmetic puzzles show that, replacing CoT-generated data with search-generated data for offline fine-tuning improves success rates by around 23% over inference-time search baselines, while reducing inference time by 180. On top of this, our learning and forgetting objective consistently outperforms both supervised fine-tuning and preference-based methods.111Code is open-source at https://github.com/twni2016/llm-reasoning-uft.
@@ 1 Introduction
Chain-of-Thought (CoT) and inference-time search (search) have been proposed to enhance LLM generalization to reasoning problems at test time. Specifically, given a task description and input , CoT (Wei et al., 2022) outputs a single reasoning path via greedy decoding, or samples multiple paths to compute average performance. In contrast, search-based methods (Lightman et al., 2023; Yao et al., 2023a; Snell et al., 2024) explicitly generate intermediate reasoning (possibly with multiple passes) before extracting via a reward model or self-evaluation. Despite inference-time search’s effectiveness, it comes with greater computational costs (e.g., Fig. 1 and Chen et al. (2024c; b)).
@@
In this paper, we propose a straightforward method consisting of three phases designed to enhance the reasoning capabilities of large language models: data generation, fine-tuning, and evaluation. First, we collect heterogeneous reasoning data from both CoT and the search-based methods, and convert the data into a unified CoT-style path format. Next, we fine-tune the base LLM by learning successful paths with the supervised fine-tuning (SFT) and forgetting unpaired failed paths with the unlikelihood loss. We refer to this objective as unlikelihood fine-tuning (UFT). Finally, we evaluate the fine-tuned model on test sets using both CoT and inference-time search, with a primary focus on efficient CoT inference. We demonstrate the effectiveness of our three-phase pipeline on two arithmetic puzzle domains: Game-of-24 and Countdown, which are widely used as testbeds for explicit search (Yao et al., 2023a; Gandhi et al., 2024). We summarize our key findings below, using Qwen2.5-Math 1.5B and 7B (Yang et al., 2024) as base LLMs:
@@ 2 Related Work
Inference-time search has shown remarkable success across reasoning (Wang et al., 2022; Stechly et al., 2024; Hao et al., 2024; Snell et al., 2024) and planning tasks (Valmeekam et al., 2023; Zheng et al., 2024; Bohnet et al., 2024). Approaches include best-of-n sampling (Nakano et al., 2021; Wang et al., 2022), multi-agent debates (Du et al., 2023), and iterative correction via predefined rules (Bai et al., 2022) or self-refinement (Shinn et al., 2023; Madaan et al., 2024). A key direction, which this paper focuses on, integrates classic search algorithms with LLMs, employing BFS (Yao et al., 2023a; Xie et al., 2023), DFS (Qin et al., 2023), A* (Zhuang et al., 2023; Meng et al., 2024a; Koh et al., 2024) and MCTS (Hao et al., 2023; Zhou et al., 2023; Zhao et al., 2024; Xie et al., 2024; Gao et al., 2024). While effective, inference-time search methods remain computationally expensive: self-refinement requires multiple expensive passes, multi-agent approaches demand substantial GPU memory, and search trees grow exponentially with depth (Chen et al., 2024c).
@@
Synthetic data, generated by classic algorithms or LLMs, is widely used for SFT in reasoning tasks to mitigate the scarcity of human-annotated datasets (Liu et al., 2024). However, despite its abundance, synthetic responses may fail to solve the task. To address this, most prior work distills reasoning knowledge to policy (Hinton et al., 2015) by selecting only correct responses. Policy distillation strategies can be categorized based on whether the source and target data originate from CoT or search algorithms. CoT-to-CoT distillation learns CoT reasoning from correct CoT paths filtered by rewards, widely used in practice (Zelikman et al., 2022; 2024; Uesato et al., 2022; Yuan et al., 2023b; Gulcehre et al., 2023; Singh et al., 2023). Search-to-search distillation either refines the proposal policy in tree search by learning from expert behavior (Feng et al., 2023; Tian et al., 2024; Zhang et al., 2024a) and preference pairs (Chen et al., 2024a), or learns a meta-policy by imitating entire search traces (Yang et al., 2022; Gandhi et al., 2024; Lehnert et al., 2024). Notably, Zhang et al. (2024c) performs search-to-CoT distillation using preference pairs from Tree-of-Thought without a verifier. In comparison, we focus on CoT, search-to-CoT distillation from diverse reasoners and distill unpaired positive and negative data to the policy, labeled by a rule-based verifier.
@@
Negative data has been extensively studied in LLM safety, where the goal is to unlearn harmful content using objectives such as Gradient Ascent and unlikelihood training (Welleck et al., 2019; Keskar et al., 2019; Yao et al., 2023b; Zhang et al., 2024b). In preference optimization, negative examples are used to increase the relative likelihood of preferred responses in paired data , commonly used in alignment (Rafailov et al., 2023; Yuan et al., 2023a; Zhao et al., 2023; Hong et al., 2024; Xu et al., 2024; Meng et al., 2024b) and reasoning (Pal et al., 2024; Pang et al., 2024; Zhang et al., 2024c; Setlur et al., 2024; Chen et al., 2024a). However, this paired format discards unpaired positives or negatives, reducing data efficiency. KTO (Ethayarajh et al., 2024) addresses the limitation of relative likelihood (Tuan & Wang, 2024) by learning from unpaired data, although it requires a reference model. In LLM reasoning, negative data is also used to train a reward model (Cobbe et al., 2021; Uesato et al., 2022; Lightman et al., 2023; Feng et al., 2023; Zhang et al., 2024a; Hosseini et al., 2024) for response re-ranking and search during inference. However, this does not inherently improve the base model as a policy. Instead of requiring preference pairs or reference models, we adopt unlikelihood training (Welleck et al., 2019) to directly forget failed paths as an auxiliary loss. Finally, concurrent work (Wang et al., 2025) highlights the challenges of unlearning in reasoning tasks, namely forgetting both incorrect reasoning traces and final answers while retaining overall reasoning ability.
@@ 4.2 Unlikelihood Fine-Tuning on Unpaired Correct and Failed Paths
While prior work typically fine-tunes LLMs on correct CoT-generated paths (Yuan et al., 2023b; Singh et al., 2023) or on preference pairs of correct and failed responses (Pang et al., 2024; Zhang et al., 2024c), we extend these approaches in two directions: (1) incorporating diverse reasoners beyond CoT to augment the training data, (2) fine-tuning the model to avoid and forget failed reasoning patterns using only failed examples. Unlike most preference-based methods, our approach leverages unpaired positive or negative data, without requiring success-failure pairs for the same question.
@@
This objective explicitly reduces the probability of incorrect paths under the model. Unlike Gradient Ascent (GA) unlearning (Yao et al., 2023b)444The GA objective is shown to be unstable to optimize (Zhang et al., 2024b; d), as it has an unbounded optimum () and a divergent gradient near the optimum. We provide results for GA in Sec. B.3 for completeness. , UL objective has a smooth optimization landscape, with gradient:
@@
Since correct paths provide explicit solutions while wrong paths only rule out alternatives, we treat UL as an auxiliary loss by setting close to 0. Moreover, as correct and failed paths may be drawn from the same search trees, they often share prefixes, introducing conflicts when is large, as also noted in preference-based methods (Zhang et al., 2024c). For example, suppose a prefix sequence () appears in both and with equal frequency. The gradient of Eq. 5 w.r.t. this prefix is , which has a stationary point at . Thus, sets the desired probability of the shared prefix, further motivating the need for a small to avoid conflicting objectives. Unlike prior work (Zhang et al., 2024c; Setlur et al., 2024; Chen et al., 2024a) that identify high-credit steps with more computation, we sidestep this with an auxiliary loss without preference pairs.
@@ 5 Experiments
We evaluate our approach on two arithmetic puzzle tasks.
@@ 5.2 Boosting Chain-of-Thought Reasoning
Incorporating high-quality training data is decisive for CoT reasoning.
@@
Lastly, we analyze the effect of forgetting loss to SFT. UFT achieves the highest mean (ignoring variance) in 26 out of 36 evaluation settings, and is statistically significantly better than baselines in 7 cases. Given the limited number of seeds (3–4), these results may underestimate the true significance, and we expect the evidence to strengthen with additional runs. In terms of magnitude, UFT yields an average improvement of 1.3% (with a maximum of 4%) in 3(b), and an average of 1.5% (with a maximum of 7%) in 2(b).
In contrast, CPO-SimPO shows mixed results. While it improves over SFT with an average gain of 1.2% and maximum of 5% in 3(b), it underperforms SFT in most scenarios in 2(b), with an average drop of 3.5%. We believe several factors contribute to the contrasting behavior of CPO-SimPO relative to SFT (and UFT) when fine-tuned on Countdown versus Game-of-24. First, as shown in Appendix Table 5, the paired preference requirement in CPO-SimPO results in the exclusion of 0.1% to 5.8% of correct data on Countdown, while no such filtering occurs for Game-of-24. Since correct data determines data quality, this reduction likely contributes to CPO-SimPO’s weaker performance on Countdown. Second, the exclusion of unpaired failed data, along with the challenge of balancing preference and NLL objectives, may have a greater impact on the more difficult Countdown task.
@@ 5.3 Mitigating Forgetting in Search-based Reasoning
6 Conclusion and Future Work
We propose a simple and effective method for improving reasoning on arithmetic puzzle domains, addressing three underexplored challenges: (1) augmenting training data by unifying mixed data formats from diverse reasoners into a CoT-style format, (2) leveraging unpaired correct and failed paths via unlikelihood fine-tuning, and (3) mitigating loss of search capability using a small learning rate. Experiments on Countdown and Game-of-24 show that data quality, unlikelihood loss, and learning rate are key to balancing CoT efficiency and search capability. Future work can explore more advanced unlearning objectives (Tamirisa et al., 2024; Huang et al., 2024), incorporate stabilization techniques (Meng et al., 2024b), and extend the framework to other reasoning-oriented LLMs.
@@ A.2 Preference-based (Reference-free) Baselines: SimPO and CPO-SimPO
Appendix B Additional Results on Countdown and Game-of-24
@@ B.1 Training Dataset Statistics
B.2 Additional Results with SFT and UFT
For completeness, we provide fine-tuning results based on Qwen2.5-Math 1.5B in Fig. 6 and Fig. 7. Given its weak search-based reasoning performance as a base model (e.g., success rate 2% in Countdown) shown in Table 6, the effects of learning rate, CoT data, and unlikelihood loss are minimal, likely due to its poor initial capability. This suggests that the observed benefits depend on the model capability.
@@ B.2 Additional Results with SFT and UFT
B.3 Additional Results with Alternative Unlearning Methods
Gradient Ascent and RepNoise.
To further evaluate our framework, we conduct additional experiments by replacing the unlikelihood (UL) loss in Eq. 5 with other established unlearning objectives, while keeping all other settings identical. In particular, we examine Gradient Ascent (GA) (Yao et al., 2023b) and RepNoise (Rosati et al., 2024).
GA is a simple baseline for unlearning that aims to minimize the probability of failed paths:
Although GA is shown to be unstable during optimization (Zhang et al., 2024b), we mitigate this issue by using a small loss coefficient of 1e-5, which is on a similar magnitude as that used for UL. This choice substantially improves numerical stability in our experiments.
Representation Noising (RepNoise) (Rosati et al., 2024), on the other hand, is a recent advanced unlearning method designed for enhancing LLM safety. It combines Gradient Ascent with removing harmful representations, encouraging the model’s representations (per-layer hidden states) for negative inputs to align with random noise. Within our framework, its objective is written as:
The first term acts as a generalized gradient-ascent objective (with an additional log-transformation), reducing the likelihood of undesired outputs based on the per-layer post-residual hidden states passed into the model’s output head , where denotes the total number of layers. The second term enforces that the pre-residual hidden states to be close to Gaussian noise according to the maximum mean discrepancy (MMD). Following the original implementation, we set the RepNoise loss coefficient to and its coefficient for the MMD term 1e-3.
CoT reasoning results for Gradient Ascent and RepNoise.
7(b) shows results for CoT reasoning with the same learning rate for all unlearning methods on Game-of-24. GA achieves performance comparable to SFT and UFT, with slightly lower averages than UFT. RepNoise, which places strong emphasis on the unlearning objective (via equal weighting and representation-level forgetting), performs consistently worse than the other methods. These observations suggest that directly applying unlearning methods originally designed for LLM safety may require additional design considerations.
Across all unlearning methods, our results consistently support the claim that higher training data quality, regardless of whether the data originate from CoT or search-based algorithms, leads to better fine-tuning performance.
Search-based reasoning results for Gradient Ascent and RepNoise.
8(b) further examines the impact of learning rate and dataset source on CoT and search-based reasoning. Following the previous setup, we sweep learning rates over 1e-6, 2e-6, 5e-6 for the 7B base model.
For GA, our key findings for SFT and UFT (Sec. 5.3) remain valid. A smaller learning rate slightly reduces CoT reasoning performance but can substantially preserve, or even improve, the base model’s search capability (e.g., 31% in ToT and 28% in RAP). Moreover, CoT training is more effective for maintaining search ability compared to mixed data (CoT+ToT+RAP), suggesting that using more on-policy data reduces distribution shift and mitigates forgetting, as discussed in Sec. 5.3.
In contrast, RepNoise exhibits a drastic drop in search-based performance (ToT and RAP), reaching 0.0% regardless of the learning rate, even though its CoT reasoning improves over the base LLM. We hypothesize that this degradation arises from the large default coefficient () in the RepNoise loss, which amplifies the effective learning rate for unlearning far beyond those used in other methods such as UFT and GA (). This aggressive unlearning, coupled with representation-level regularization toward random noise, likely disrupts internal representations crucial for search behaviors (e.g., hypothesis branching and reward estimation), even when the training objective (CoT reasoning) itself remains stable. These results underscore the importance of carefully tuning the learning rate and loss coefficient when extending unlearning objectives to LLM reasoning domains.
| SFT | UFT | GA | RepNoise | |
| CoT data (quality: 92.9%) | ||||
| succ (greedy) | ||||
| succ (pass@1) | ||||
| ToT+RAP data (quality: 69.3%) | ||||
| succ (greedy) | ||||
| succ (pass@1) | ||||
| CoT+ToT+RAP data (quality: 95.3%) | ||||
| succ (greedy) | ||||
| succ (pass@1) |
| SFT | UFT | GA | RepNoise | |
| CoT data (quality: 82.1%) | ||||
| succ (greedy) | ||||
| succ (pass@1) | ||||
| ToT+RAP data (quality: 44.7%) | ||||
| succ (greedy) | ||||
| succ (pass@1) | ||||
| CoT+ToT+RAP data (quality: 87.6%) | ||||
| succ (greedy) | ||||
| succ (pass@1) |
| Learning rate | CoT (greedy) | ToT | RAP | |
| Base LLM (Qwen2.5-Math 7B) | N/A | 6.0% | 28.0% | 27.0% |
| CoT data (quality: 92.9%) | 5e-6 | |||
| 2e-6 | ||||
| 1e-6 | ||||
| ToT+RAP data (quality: 69.3%) | 5e-6 | |||
| 2e-6 | ||||
| 1e-6 | ||||
| CoT+ToT+RAP data (quality: 95.3%) | 5e-6 | |||
| 2e-6 | ||||
| 1e-6 |
| Learning rate | CoT (greedy) | ToT | RAP | |
| Base LLM (Qwen2.5-Math 7B) | N/A | 6.0% | 28.0% | 27.0% |
| CoT data (quality: 92.9%) | 5e-6 | |||
| 2e-6 | ||||
| 1e-6 | ||||
| ToT+RAP data (quality: 69.3%) | 5e-6 | |||
| 2e-6 | ||||
| 1e-6 | ||||
| CoT+ToT+RAP data (quality: 95.3%) | 5e-6 | |||
| 2e-6 | ||||
| 1e-6 |